Is the BOLD and/or UNDERLINE CORRECT:

Cho currently earns a nominal wage of $12.00 per hour; in other words, the amount of her paycheck each week is $12.00 per hour times the number of hours she works. Suppose the price of orange juice is $2.40 per gallon; in this case, Cho's real wage, in terms of the amount of orange juice she can buy with her paycheck, is 5 gallons of orange juice per hour.

When workers and firms negotiate compensation packages, they have expectations about the price level (and changes in the price level) and agree on a nominal   wage with those expectations in mind. If the price level turns out to be lower than expected, a worker's real wage is lower than both the worker and employer expected when they agreed to the wage.

Cho and her employer both expected inflation to be 4% between 2012 and 2013, so they agreed, in a two-year contract, that she would earn $12.00 per hour in 2012 and $12.48 per hour in 2013. However, suppose inflation between 2012 and 2013 actually turned out to be 2%, not 4%. For example, suppose the price of orange juice rose from $2.40 per gallon to $2.45 per gallon. This means that between 2012 and 2013, Cho's nominal wage (increase/decrease) by, ____% and her real wage   (increase/decrease) by approximately _____% .
Answer #1

a) If the price level turns out to be lower than expected, a worker's real wage is HIGHER than both the worker and employer expected when they agreed to the wage.

Reason : Real wage = Nominal wage / Price

Lower price (lower denominator) => Higher real wage

b) % increase in nominal wage : [(12.48 - 12) / 12] x 100 = 4%

c) Real wage (@ inflation = 4%) = 12/4 = $3 (because they expected inflation to be 4%)

Real wage (@ inflation = 2%) = 12.48/2 = $6.24 (because the actual inflation turned out to be 2%)

% increase in real wage : [(6.24 - 3) / 3] x 100 = 108%